what is the integrating factor of the non-exact differential equation (3x^2+y^3+4)dx+(x^2-2xy)dy=0


1
Expert's answer
2021-03-22T03:26:52-0400

"(3x^2+y^3+4)dx+(x^2-2xy)dy=0\\
M=3x^2+y^3+4, N=x^2-2xy\\
\frac{dM}{dy}=3y^2, \frac{dN}{dx}=2x-2y\\
\frac{dM}{dy} \ne \frac{dN}{dx}\\
\text{multiply through by}, x^my^n\\
x^my^n(3x^2+y^3+4)dx+ x^my^n(x^2-2xy)dy=0\\
(3x^{m+2}y^n+x^my^{n+3}+4x^my^n)dx+(x^{m+2}y^n-2x^{m+1}y^{n+1})dy=0\\
M=3x^{m+2}y^n+x^my^{n+3}+4x^my^n\\
N=x^{m+2}y^n-2x^{m+1}y^{n+1}\\
\frac{dM}{dy}=3nx^{m+2}y^{n-1}+(n+3)x^my^{n+2}+4nx^my^{n-1}\\
\frac{dN}{dx}=(m+2)x^{m+1}y^n-2(m+1)x^my^{n+1}\\
\frac{dM}{dy}=\frac{dN}{dx}\\
3nx^{m+2}y^{n-1}+(n+3)x^my^{n+2}+4nx^my^{n-1}=(m+2)x^{m+1}y^n-2(m+1)x^my^{n+1}\\
3nx^{m+2}y^{n-1}+(n+3)x^my^{n+2}+4nx^my^{n-1}-(m+2)x^{m+1}y^n+2(m+1)x^my^{n+1}=0\\
\text{set the coeefficients to be 0, to get the value of m and n}\\
\text{integrating factor is}\\, x^my^n \\
\text{for the value of m, n}\\
3n=0, n=0\\
n+3=0, n=-3\\
4n=0, n=0\\
-m-2=0, m=-2\\
2m+2=0, m=-1\\
\text{the integrating factor becomes}\\
x^0y^{-2}"